# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v3

@@ 3.2 Deterministic equivalents of the diagonal of the resolvent
Let $Q_{p}(z)=(\widehat{\Sigma}_{n}-zI_{p})^{-1}$ for $z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$ be the resolvent of the matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$, and for any square matrix $A$, we denote by $\Delta[A]$ the diagonal matrix whose diagonal entries are those of $A$.
Then, as shown in the next subsection, a key argument to derive deterministic equivalents of the predictive and training risks is to prove that the matrix $T_{p}(z)$, defined by (2.3), is a relevant deterministic equivalent of the diagonal matrix $\Delta[Q_{p}(z)]$ for an appropriate notion of asymptotic equivalence between matrices of growing size. This is the purpose of Theorem 3.1 below that derives a stronger convergence result than the one stated in Proposition 2.1 which only shows that $\frac{1}{p}\operatorname{Tr}[T_{p}(z)]$ is a deterministic equivalent of $g_{\hat{\mu}_{n}}(z)=\frac{1}{p}\operatorname{Tr}[Q_{p}(z)]$. To this end, we define the following equivalence relation in order to specify an appropriate notion of deterministic equivalent for the matrix $\Delta[Q_{p}(-\lambda)]$.
@@
Then using this definition, we extend Theorem 2.1 from [HLN07] by proving that $T_{p}(z)$ is a relevant deterministic equivalent of $\Delta[Q_{p}(z)]$ in the following sense.
@@
Define the following matrix families $\Delta[\mathbf{Q}(z)]=(\Delta[Q_{p}(z)])_{p\geq 1}$, $\Delta[\mathbf{Q^{\prime}}(z)]=(\Delta[Q^{\prime}_{p}(z)])_{p\geq 1}$, $\mathbf{T}(z)=(T_{p}(z))_{p\geq 1}$ and $\mathbf{T^{\prime}}(z)=(T^{\prime}_{p}(z))_{p\geq 1}$. Then, under Assumptions 1.2-1.3 the following equivalences hold:
@@
We postpone the proof of this theorem to Appendix A.2 that is inspired from the proof of Theorem 2.1 in [HLN07]. This theorem is a cornerstone in the derivation of deterministic equivalents for $\hat{r}_{\lambda}^{test}(X_{n})$ and $\hat{r}_{\lambda}^{train}(X_{n})$ that leads to the main results of this paper provided in the next subsection.
3.3 Deterministic equivalents of the training and predictive risks
@@
$\mathop{\rm Bias}\nolimits(\hat{\theta}_{\lambda})$ $=$ $\mathbb{E}\big{[}(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}]-\beta_{\ast})^{\top}\widetilde{S}_{p}(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}]-\beta_{\ast})\big{|}X_{n}\big{]}$
$=$ $\frac{\lambda^{2}\alpha^{2}}{p}\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big{]},$
and
$\mathop{\rm Var}(\hat{\theta}_{\lambda})$ $=$ $\mathbb{E}\big{[}(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}])^{\top}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}])\big{|}X_{n}\big{]}$
$=$ $\frac{\sigma^{2}}{n}\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}(-\lambda)]-\lambda\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big{]}.$
We can now give a deterministic equivalent of the predictive risk $\hat{r}_{\lambda}^{test}(X_{n})$ that is obtained in a simple way by replacing the diagonal matrix $\Delta[Q_{p}(z)]$ with the deterministic matrix $T_{p}(z)$ in the expression (3.5) of the predictive risk. All proofs of the following results are given in Appendix A.3.
@@
Under Assumptions 1.1 to 1.3, one can provide a deterministic equivalent for the predictive risk (respectively for the training risk), denoted by $r_{\lambda}^{test}$ (respectively $r_{\lambda}^{train}$) and defined as follows
@@
in the sense that it satisfies
$\lim_{n\to\infty,\;p/n\to c}|\hat{r}_{\lambda}^{\bullet}(X_{n})-r_{\lambda}^{\bullet}(X_{n})|=0,\;\mbox{almost surely},$
@@
where $\bullet\in\{train,\ test\}$ and $T_{p}^{\prime}(z)$ denotes the derivative of $T_{p}(z)$ with respect to $z$.
@@
From Theorem 3.2, one can deduce that, in the high dimensional regime, the training and predictive risk crystallize around deterministic values that depend on the data only through their variance profile. These deterministic equivalents can be expressed explicitly whenever the variance profile is quasi doubly-stochastic since $T_{p}(z)$ is equal to the Marchenko-Pastur Stieltjes transform in this case (see Section 2.3). However, the equivalents are not explicit in the general case since one does not have an explicit formula for $T_{p}(z)$. Nevertheless, $T_{p}(z)$ being the solution of a fixed-point equation, it can be approximated through a fixed-point algoritheorem. This alternative allowed us to perform numerical experiment that testify the accuracy of these deterministic equivalents (see Section 4).
@@
Under Assumptions 1.2 to 1.5, if $\frac{p}{n}<1$ then $T_{p}(-\lambda)$ and its derivative, $T_{p}^{\prime}(-\lambda)$ admit a limit when $\lambda$ tends to 00. Indeed, there exists a constant $\tau>0$ such that
@@
On the other hand, if $\frac{p}{n}>1$ then $\kappa(\lambda)$ and its derivative $\kappa^{\prime}(\lambda)$ also admit a limit when $\lambda$ tends to 00 and we denote these limits by
@@
This lemma asserts that if $p<n$ (respectively $p>n$) then $T_{p}(-\lambda)$ and $T^{\prime}_{p}(-\lambda)$ (respectively $\kappa(\lambda)$) admit limits whenever $\lambda$ goes to 00. These limits ensures that the deterministic equivalents provided by Theorem 3.2 exist in the ridge (less) case, that is $\lambda=0$ (See Corollary 3.2).
@@
Suppose Assumptions 1.2-1.5 and that the variance profile $\Gamma_{n}$ is quasi doubly stochastic. Then, if $\frac{p}{n}<1$, one has that $T_{p}(0^{-})=m_{p}(0)I_{p},$ and, if $\frac{p}{n}>1$, $\kappa(0^{+})=\frac{1}{\tilde{m}_{n}(0)}I_{p}\quad\mathrm{and}\quad\kappa^{\prime}(0^{+})=\frac{\tilde{m}_{n}^{\prime}(0)}{\tilde{m}_{n}^{2}(0)}I_{p},$ where $m_{p}(.)$, resp. $\tilde{m}_{n}(.)$, is the Stieltjes transform of the Marchenko-Pastur distribution with parameter $p/n$, resp. $n/p$.
Then, Theorem 3.2 allows us to understand the behavior of $r_{\lambda}^{test}(X_{n})$ when $\lambda\rightarrow 0$ and $\lambda\rightarrow+\infty$ through the following corollary.
@@
Under Assumptions 1.1 to 1.3, as $\lambda\rightarrow+\infty$, the limit of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ of the predictive risk is
$\lim_{\lambda\rightarrow+\infty}r_{\lambda}^{test}(X_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}.$
Moreover, under Assumptions 1.1 to 1.5, as $\lambda\rightarrow 0$, the limit of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ is as follows
if $\frac{p}{n}<1$ then
$\lim_{\lambda\rightarrow 0^{+}}r_{\lambda}^{test}(X_{n})=\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(0^{-})]+\sigma^{2},$
if $\frac{p}{n}>1$ then
$\lim_{\lambda\rightarrow 0^{+}}r_{\lambda}^{test}$ $=$ $\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}\kappa(0^{+})(\Sigma_{n}+\kappa(0^{+}))^{-1}]$
$+\frac{\sigma^{2}}{n}\mathrm{Tr}[\kappa^{\prime}(0^{+})\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(0^{+}))^{-2}]+\sigma^{2},$
@@
In the case $\lambda\rightarrow 0$, Corollary 3.2 yields a deterministic equivalent of the predicitve risk of the mininum norm least-square estimator $\hat{\theta}$. Then, it can be seen that the risk $r_{0}(X_{n})$ exhibit two different behaviors depending on the value of the ratio $p/n$ with respect to one.
@@
it follows that
$r_{0}(X_{n})=\frac{\sigma^{2}}{1-\frac{p}{n}}\times\frac{1}{n}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}=\sigma^{2}\frac{\frac{p}{n}}{1-\frac{p}{n}}+\sigma^{2}\rightarrow\sigma^{2}\frac{c}{1-c}+\sigma^{2},$
as $n\to\infty,\;\frac{p}{n}\to c$, which corresponds, when $c<1$, to the known asymptotic limit of the predictive risk of least squares estimation for iid data when the entries of the features matrix $X_{n}$ are iid centered random variables with variances equal to $1$, see e.g. [HMRT22][Proposition 2].
@@
Hence, using that $\frac{1}{n}\mathrm{Tr}[\widetilde{S}_{p}]=p/n$, we finally obtain that
$r_{0}(X_{n})=\alpha^{2}\left(1-\frac{n}{p}\right)+\sigma^{2}\frac{1}{p/n-1}\rightarrow\alpha^{2}\left(1-\frac{1}{c}\right)+\sigma^{2}\frac{1}{c-1},$
as $n\to\infty,\;\frac{p}{n}\to c$, which corresponds, when $c>1$, to known results on the bias-variance decomposition of the asymptotic limit of the predictive risk of the minimum norm least squares estimator for iid data when the entries of $X_{n}$ are iid centered random variables with variance $1$, see e.g. [HMRT22][Theorem 1].
@@
Nevertheless, in Section 4, we use Corollary 3.2 and computational methods to evaluate $T_{p}(0^{-})$ and $\kappa(0^{+})$ to report numerical experiments illustrating that the double descent phenomemum for the predictive risk of $\hat{\theta}$ also holds in the high-dimensional model (1.1) with more general variance profile than a quasi-bistochastic one. In Section 4, we also exhibit variance profiles for which the predictive risk has a shape that differs from double descent.
Corollary 3.3.
Under Assumptions 1.1 to 1.3, the function $\lambda\mapsto\hat{r}_{\lambda}^{test}(X_{n})$ reaches its minimium at $\lambda_{\ast}=\frac{\sigma^{2}p}{\alpha^{2}n}$ meaning that $\hat{r}_{\lambda_{\ast}}^{test}(X_{n})\leq\hat{r}_{\lambda}^{test}(X_{n})$, for any $\lambda>0$.
Note that although $\hat{r}_{\lambda}^{test}(X_{n})$ depends on the variance profile, this is not the case of the optimal value of $\lambda_{\ast}$ that minimizes the predictive risk as shown by Corollary 3.3. Moreover, the optimal value $\lambda_{\ast}$ is also the one minimizing $\lambda\mapsto\hat{r}_{\lambda}^{test}(X_{n})$ in the framework of [DW18] with a matrix $X_{n}$ of features made of iid rows.

### v4

@@ 3.2 Deterministic equivalents of the diagonal of the resolvent
Let $Q_{p}(z)=(\widehat{\Sigma}_{n}-zI_{p})^{-1}$ for $z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$ be the resolvent of the matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$, and for any square matrix $A$, we denote by $\Delta[A]$ the diagonal matrix whose diagonal entries are those of $A$. Then, as shown in the next subsection, a key argument to derive deterministic equivalents of the predictive and training risks is to prove that the matrix $T_{p}(z)$, defined by Equation (2.2), is a relevant deterministic equivalent of the diagonal matrix $\Delta[Q_{p}(z)]$ for an appropriate notion of asymptotic equivalence between matrices of growing size. This is the purpose of Theorem 3.1 below that derives a stronger convergence result than the one stated in Proposition 2.1 which only shows that $\frac{1}{p}\operatorname{Tr}[T_{p}(z)]$ is a deterministic equivalent of $g_{\hat{\mu}_{n}}(z)=\frac{1}{p}\operatorname{Tr}[Q_{p}(z)]$. To this end, we define the following equivalence relation in order to specify an appropriate notion of deterministic equivalent for the matrix $\Delta[Q_{p}(-\lambda)]$.
@@
Then using this definition, we extend Proposition 2.1 from [HLN07] by proving that $T_{p}(z)$ is a relevant deterministic equivalent of $\Delta[Q_{p}(z)]$ in the following sense.
@@
Define the following matrix families $\Delta[\mathbf{Q}(z)]=(\Delta[Q_{p}(z)])_{p\geq 1}$, $\Delta[\mathbf{Q^{\prime}}(z)]=(\Delta[Q^{\prime}_{p}(z)])_{p\geq 1}$, $\mathbf{T}(z)=(T_{p}(z))_{p\geq 1}$ and $\mathbf{T^{\prime}}(z)=(T^{\prime}_{p}(z))_{p\geq 1}$, where $Q_{p}^{\prime}(z)$ and $T_{p}^{\prime}(z)$ are the derivative of $Q_{p}(z)$ and $T_{p}(z)$ with respect to $z$. Then, supposing that Assumptions 1.2 and 1.3 hold, one has the following equivalences:
@@
We postpone the proof of this theorem to Section A.2 that is inspired from the proof of Proposition 2.1 in [HLN07]. This theorem is a cornerstone in the derivation of deterministic equivalents for $\hat{r}_{\lambda}^{test}(X_{n})$ and $\hat{r}_{\lambda}^{train}(X_{n})$ that leads to the main results of this paper provided in the next subsection.
3.3 Deterministic equivalents of the training and predictive risks for $\lambda>0$
@@
$\mathop{\rm Bias}\nolimits(\hat{\theta}_{\lambda})=\mathbb{E}\big[(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}]-\beta_{\ast})^{\top}\widetilde{S}_{p}(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}]-\beta_{\ast})\big|X_{n}\big]=\frac{\lambda^{2}\alpha^{2}}{p}\mathrm{Tr}\big[\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big],$
$\mathop{\rm Var}(\hat{\theta}_{\lambda})=\mathbb{E}\big[(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}])^{\top}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}])\big|X_{n}\big]=\frac{\sigma^{2}}{n}\mathrm{Tr}\big[\widetilde{S}_{p}\Delta[Q_{p}(-\lambda)]-\lambda\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big].$
We can now give a deterministic equivalent of the predictive risk $\hat{r}_{\lambda}^{test}(X_{n})$ that is obtained in a simple way by replacing the diagonal matrix $\Delta[Q_{p}(z)]$ with the deterministic matrix $T_{p}(z)$ in the expression Equation (3.5) of the predictive risk. All proofs of the following results are given in Section A.3.
@@
Suppose that Assumption 1.1 to Assumption 1.3 hold true, then one can provide a deterministic equivalent for the predictive risk (respectively for the training risk), denoted by $r_{\lambda}^{test}$ (respectively $r_{\lambda}^{train}$) and defined as follows
@@
in the sense that it satisfies $\lim_{n\to\infty,\;p/n\to c}|\hat{r}_{\lambda}^{\bullet}(X_{n})-r_{\lambda}^{\bullet}(X_{n})|=0,\;\mbox{almost surely},$ where $\bullet\in\{train,\ test\}$ and $T_{p}^{\prime}(z)$ denotes the derivative of $T_{p}(z)$ with respect to $z$.
@@
From Theorem 3.2, one can deduce that, in the high dimensional regime, the training and predictive risk crystallize around deterministic values that depend on the data only through their variance profile. These deterministic equivalents can be expressed explicitly whenever the variance profile is quasi doubly-stochastic since $T_{p}(z)$ is equal to the Marchenko-Pastur Stieltjes transform in this case (see Section 2.3). However, the equivalents are not explicit in the general case since one does not have an explicit formula for $T_{p}(z)$. Nevertheless, $T_{p}(z)$ being the solution of a fixed-point equation, it can be approximated through a fixed-point algorithm. This alternative allowed us to perform numerical experiment that testify the accuracy of these deterministic equivalents (see Section 4).
We now study the choice of an optimal regularization parameter $\lambda$ using either the empirical predictive risk or its deterministic equivalent.
Corollary 3.1.
Suppose that Assumption 1.1 to Assumption 1.3 hold true. Then, both functions $\lambda\mapsto\hat{r}_{\lambda}^{test}(X_{n})$ and $\lambda\mapsto r_{\lambda}^{test}(X_{n})$ reach their minimum at $\lambda_{\ast}=\frac{\sigma^{2}p}{\alpha^{2}n}$ meaning that $\hat{r}_{\lambda_{\ast}}^{test}(X_{n})\leq\hat{r}_{\lambda}^{test}(X_{n})$ and $r_{\lambda_{\ast}}^{test}(X_{n})\leq r_{\lambda}^{test}(X_{n})$, for any $\lambda>0$.
Note that although the expression of $r_{\lambda}^{test}(X_{n})$ depends on the variance profile, this is not the case of the optimal value of $\lambda_{\ast}$ that minimizes the predictive risk as shown by Corollary 3.1. Moreover, the optimal value $\lambda_{\ast}$ is also the one minimizing the predictive risk in the framework of [DW18] with a matrix $X_{n}$ of features made of iid rows. Hence, Corollary 3.1 shows that an optimal regularization parameter derived in the standard iid setting remains valid under the more general non-iid framework using a variance profile, thereby reinforcing the universality of the optimal value $\lambda_{\ast}$ for predictive risk minimization. Interestingly, both empirical and deterministic predictive risk achieve their minimum at the same optimal regularization parameter. Nevertheless, this strategy leads to the selection of $\lambda$ that requires to have access to test data or prior knowledge of the signal strength $\alpha$.
Building upon the deterministic equivalents derived above and the use of the GCV criteria (1.6), one may also adopt the following alternative strategy for selecting the regularization parameter $\lambda$. Introducing the following deterministic equivalent
$\color[rgb]{0,0,0}G(\lambda)=\frac{r_{\lambda}^{\text{train}}(X_{n})}{\left(1-df_{1}(\lambda)\right)^{2}},$}
one obtains that
$\lim_{n\to\infty,\;p/n\to c}\big|\hat{G}(\lambda)-G(\lambda)\big|=0\qquad\text{almost surely}.$
Hence, in the proportional asymptotic regime, the empirical GCV criteria concentrates around a deterministic limit that depends on the variance profile only through the deterministic equivalents of the training risk and the degrees of freedom.
Figure 2: Comparison of the predictive risk, the GCV and their deterministic equivalents for several variances profiles: (a) block variance profile, (b) alternated columns variance profile, (c) polynomial variance profile, as described in Section 4 on numerical experiments, with $\lambda$ ranging from $0.1$ to $3$, $\alpha=1$, $\sigma=1$, $n=1000$ and $p=500$.
This naturally leads to the selection of a regularization parameter defined as
$\lambda^{\mathrm{GCV}}\in\arg\min_{\lambda>0}G(\lambda).$ (3.8)
However, this estimator is not guaranteed to recover the oracle parameter $\lambda_{\ast}$ derived in Corollary 3.1 as illustrated in Figure 2, where the predictive risk and the GCV criterion attain their minima at distinct, albeit nearby, values of $\lambda$. Hence, while the oracle parameter $\lambda_{\ast}$ is universal and does not depend on the variance profile, the GCV-selected parameter $\lambda^{\mathrm{GCV}}$ depends on it.
Since $G(\lambda)$ is defined in terms of the deterministic equivalent $\widetilde{T}(-\lambda)$, no explicit closed-form expression for $G(\lambda)$ is available in the general case. When the variance profile is quasi doubly stochastic, one has that $T_{p}(z)=m_{p}(z)I_{p}$ and $df_{1}(\lambda)=\left(1+\frac{1}{\tilde{m}_{n}(-\lambda)}\right)^{-1},$ where $m_{p}(\cdot)$ and $\tilde{m}_{n}(\cdot)$ are Stieltjes transforms related to the Marchenko-Pastur distribution satisfying (2.3) and (2.4). Consequently, under the assumption of a quasi doubly stochastic variance profile, we thus obtain that
$G(\lambda)=\frac{1+\tilde{m}_{n}(-\lambda)}{\tilde{m}_{n}(-\lambda)}\left(\frac{\lambda^{2}\alpha^{2}p}{n}\left(m_{p}(-\lambda)-\lambda m_{p}^{\prime}(-\lambda)\right)+\frac{\lambda^{2}\sigma^{2}p}{n}m_{p}^{\prime}(-\lambda)\right),$
but finding a closed-form minimizer of this functional remains unclear.
3.4 Deterministic equivalents of the training and predictive risks for $\lambda=0$
Note that the previous results hold for $\lambda>0$, however one can extends the study for $\lambda=0$, by considering the following additional hypotheses :
Assumption 3.1.
Assumption 3.2.
It is proved in [AEK17b] that Assumptions 3.1 and 3.2 ensure that the spectrum of $\frac{1}{n}X_{n}^{\top}X_{n}$ is bounded away from 0, which is a necessary condition to study the $\lambda=0$ case. Moreover, this case is of interest since it gives birth to the double descent phenomenon described in Section 4.
@@ 3.4 Deterministic equivalents of the training and predictive risks for $\lambda=0$
Suppose that Assumption 1.2 to Assumption 3.2 hold true, if $\frac{p}{n}<1$ then $T_{p}(-\lambda)$ and its derivative, $T_{p}^{\prime}(-\lambda)$ admit a limit when $\lambda$ tends to 0. Indeed, there exists a constant $\tau>0$ such that
@@
On the other hand, if $\frac{p}{n}>1$ then $\kappa(\lambda)$ and its derivative $\kappa^{\prime}(\lambda)$ also admit a limit when $\lambda$ tends to 0 and we denote these limits by
@@
This lemma asserts that if $p<n$ (respectively $p>n$) then $T_{p}(-\lambda)$ and $T^{\prime}_{p}(-\lambda)$ (respectively $\kappa(\lambda)$) admit limits whenever $\lambda$ goes to 0. These limits ensures that the deterministic equivalents provided by Theorem 3.2 exist in the ridge (less) case, that is $\lambda=0$ (See Corollary 3.3).
@@
Suppose that Assumption 1.2 to Assumption 3.2 hold true and that the variance profile $\Gamma_{n}$ is quasi doubly stochastic. Then,
    if $\frac{p}{n}<1$, one has that $T_{p}(0^{-})=m_{p}(0)I_{p},$
    if $\frac{p}{n}>1$, $\kappa(0^{+})=\frac{1}{\tilde{m}_{n}(0)}I_{p}\quad\mathrm{and}\quad\kappa^{\prime}(0^{+})=\frac{\tilde{m}_{n}^{\prime}(0)}{\tilde{m}_{n}^{2}(0)}I_{p},$
where $m_{p}(.)$, resp. $\tilde{m}_{n}(.)$, is the Stieltjes transform of the Marchenko-Pastur distribution with parameter $p/n$, resp. $n/p$.
The deterministic equivalents provided in Theorem 3.2 heavily depend on $T_{p}(-\lambda)$ and $\kappa(\lambda)$ and Corollary 3.2 ensures that these values can be evaluated for $\lambda=0$ as long as $p\neq 0$. Hence, one can understand the behavior of $r_{\lambda}^{test}(X_{n})$ when $\lambda\rightarrow 0$, that depends on the ratio $\frac{p}{n}$. The following corollary exploits Corollary 3.2 in order to describe this behavior that gives birth to the double descent phenomenon as we will see in Section 4.
@@
Suppose that Assumption 1.1 to Assumption 1.3 hold true, then the limit of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ of the predictive risk as $\lambda\rightarrow+\infty$ is
$\lim_{\lambda\rightarrow+\infty}r_{\lambda}^{test}(X_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2},$
Moreover, supposing that Assumption 1.1 to Assumption 3.2 hold true, then the limit of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ , as $\lambda\rightarrow 0$, is as follows
    if $\frac{p}{n}<1$ then $\lim_{\lambda\rightarrow 0^{+}}r_{\lambda}^{test}(X_{n})=\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(0^{-})]+\sigma^{2},$
    if $\frac{p}{n}>1$ then $\lim_{\lambda\rightarrow 0^{+}}r_{\lambda}^{test}=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}\kappa(0^{+})(\Sigma_{n}+\kappa(0^{+}))^{-1}]+\frac{\sigma^{2}}{n}\mathrm{Tr}[\kappa^{\prime}(0^{+})\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(0^{+}))^{-2}]+\sigma^{2},$
@@
In the case $\lambda\rightarrow 0$, Corollary 3.3 yields a deterministic equivalent of the predictive risk of the minimum norm least-square estimator $\hat{\theta}$. Then, it can be seen that the risk $r_{0}^{test}(X_{n})$ exhibit two different behaviors depending on the value of the ratio $p/n$ with respect to one. Moreover, it can be seen in Section 4 that $r_{0}^{test}(X_{n})$ seems to be an accurate estimation of the actual predictive risk in the ridgeless case, namely $\hat{r}_{0}^{test}(X_{n})$.
@@
it follows that $r_{0}(X_{n})=\frac{\sigma^{2}}{1-\frac{p}{n}}\times\frac{1}{n}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}=\sigma^{2}\frac{\frac{p}{n}}{1-\frac{p}{n}}+\sigma^{2}\rightarrow\sigma^{2}\frac{c}{1-c}+\sigma^{2},$ as $n\to\infty,\;\frac{p}{n}\to c$, which corresponds, when $c<1$, to the known asymptotic limit of the predictive risk of least squares estimation for iid data when the entries of the features matrix $X_{n}$ are iid centered random variables with variances equal to $1$, see e.g. [HMRT22][Proposition 2].
@@
Hence, using that $\frac{1}{n}\mathrm{Tr}[\widetilde{S}_{p}]=p/n$, we finally obtain that $r_{0}(X_{n})=\alpha^{2}\left(1-\frac{n}{p}\right)+\sigma^{2}\frac{1}{p/n-1}\rightarrow\alpha^{2}\left(1-\frac{1}{c}\right)+\sigma^{2}\frac{1}{c-1},$ as $n\to\infty,\;\frac{p}{n}\to c$, which corresponds, when $c>1$, to known results on the bias-variance decomposition of the asymptotic limit of the predictive risk of the minimum norm least squares estimator for iid data when the entries of $X_{n}$ are iid centered random variables with variance $1$, see e.g. [HMRT22][Theorem 1].
@@
Nevertheless, in Section 4, we use Corollary 3.3 and computational methods to evaluate $T_{p}(0^{-})$ and $\kappa(0^{+})$ to report numerical experiments illustrating that the double descent phenomenon for the predictive risk of $\hat{\theta}$ also holds in the high-dimensional model Equation (1.2) with more general variance profile than a quasi-bistochastic one. In Section 4, we also exhibit variance profiles for which the predictive risk has a shape that differs from double descent.